The total electric motor of a model train accelerates the train from rest to 0.620 m/s in 21.0cs the total mass of the train is 875g find the average power delivered to the train during the accelerates


1
Expert's answer
2022-04-28T16:42:35-0400
"P=\frac{mv^2}{2t}=\frac{0.875*0.620^2}{2*21.0*10^{-3}}=8.0\:\rm W"